OpenStudy (anonymous):

Ellen is making a jacket, which costs $42 at the store. She will need 5 yd. of fabric at $4.79 per yard, a pattern that costs $2.50, and a zipper that costs $3.15. She has thread at home. About how much will she save by making the jacket herself

OpenStudy (anonymous):

well you would first multiply 4.79* 5 =23.95 then add 2.50 to that (23.95+2.5 =26.45) then add 3.15 (26.45+3.15=29.60) subtract that total from the original cost (42-29.60=12.40) she saves $12.40

OpenStudy (anonymous):

thank yu